# Embodied Crowd Counting

### v1

@@ 3.3.1 Overview
Previous embodied navigation agents are designed for indoor environments [27], or relying on language assistance [25, 45], making them restricted in large-scale outdoor scenes. Under such context, a zero-shot agent, ZECC, which can actively control altitude and conduct crowd analysis, is proposed. As illustrated in Figure 2, the method consists of 2 components, Active Top-down Exploration (ATE) and Normal-line based Navigation (NLBN). ATE is designed for adjusting agent altitude for efficient coarse crowd distribution estimation, and NLBN estimates normal lines on top of crowd for accurate crowd observation, alleviating occlusion.
@@ 3.3.2 Active Top-down Exploration (ATE)
Crowds are often concentrated in specific areas, such as roads and squares, making it unnecessary to explore every region of the environment. To address this, ATE is proposed to estimate the global distribution of crowds from a high altitude. This approach aims to improve efficiency by focusing exploration efforts on the most relevant areas and avoiding unnecessary locations. In particular. high-altitude exploration (HAE) brings a broader field of view for decision making, as well as relatively less exploration cost since obstacles are often sparse in high altitude. While low-altitude exploration (LAE) provides close range observation for clear target detection. It is essential for the agent to plan HAE and LAE to achieve efficiency and accuracy simultaneously. Therefore, ATE leverages the Z-axis mobility of outdoor agents and the common sense of multi-modal large language model (MLLM) to switch between HAE and LAE by local environment layout reasoning. Then, the crowd distribution is estimated by crowd counting model to predict density maps in the environment.
Specifically, the agent collects observations ${O_{t}}=\left\{{o_{t}^{1},\ldots,o_{t}^{c}}\right\}$ and pose $p_{t}$ at time step $t$, where $c$ indicates the camera number of the agent. Then, a MLLM is prompted by using the observations and text prompts to conduct environment layout reasoning. The MLLM is asked to predict whether current location is valuable for LAE, by refering current crowd appearence and obstacle layout. This process is formulate as:
@@
where $\text{MLLM}\left(\cdot\right)$ is the inference process and $I$ is the prompt. $s_{t}\in\left[{0,1}\right]$. If $s_{t}>0.5$, the agent will adjust its altitude for LAE. After switching altitude strategy, the agent will conduct regular exploration. For HAE, one of the frontiers between explored and unknown areas is selected as navigation point. For LAE, the agent will keep exploring frontiers until areas within its field of view during HAE are fully explored. During LAE, once the agent reaches a frontier $f$, it gathers observations ${O_{f}}=\left\{{o_{f}^{1},\ldots,o_{f}^{c}}\right\}$, and a crowd counting model predicts crowd density maps on the observations. Then the density maps are projected onto the global point cloud to form a global crowd distribution. This is formulated as:
@@ 3.3.3 Normal-line based Navigation (NLBN)
NLBN is utilized for the precise analysis of potential overlapping structures within crowds. The intuition behind this approach is that while a random viewpoint may often obscure visibility in a dense crowd, there are several vantage points that can still provide a clear view for identifying individuals. These observation points are expected to be above and maintain a certain degree to the center of the crowd, since a top-down view can distinguish overlapping, and perspectives from certain angles can ensure that the targets are in the field of view (FOV) of the agent. To find such observation points, NLBN transfers the crowd detection task into a surface detection task. By fitting surfaces on the crowd distribution, normal lines of the surfaces can be obtained, and view points which fulfill the constraints mentioned above can be accurately calculated by using normal lines.
After achieving global crowd distribution in ATE, NLBN first clusters the crowd point cloud into subregions using a clustering method. This step helps divide the large crowd point cloud into manageable parts, as navigating to each navigation point can be resource-intensive. Next, surfaces are fitted to these subregions based on the point cloud data to obtain normal lines. This process transforms the challenging task of crowd analysis into a more manageable surface analysis. Finally, optimized navigation points are generated based on the normal lines, ensuring they maintain a certain distance and angle from the crowd surfaces for accurate crowd detection. A viewpoint-based method is employed to confirm that the selected navigation points do not overlap with any obstacles.
In particular, GMM [35] is used to divide the global crowd distribution into manageable subregions. It ensures each sample point in a distribution can be assigned with a cluster. Thus, it can divide any crowd distribution into patches. A parameter $\epsilon$ is used to determine the size of each GMM cluster.
After the clusters and centers are generated, surfaces are fitted for each patch. Then, for the $N$-th cluster, the normal is obtained and represented by ${\bf{d}}_{N}^{{\rm{cluster}}}$. Candidate view directions $\{\mathbf{d}_{N1}^{\text{view}},...,\mathbf{d}_{Nm}^{\text{view}}\}_{N}$ are sampled under angular constraints:
@@
where $\zeta$ is a hyper parameter. These candidate view directions can bring the agent to a optimized observation point by selecting a position along the vector. However, there are two issues: 1) The position may be located on obstacles; 2) The crowd cluster may be out of the agent’s FOV. To solve the issues and identify the final navigation points from the potential ones, a simple yet effective technique is proposed. In the process of ATE, each cluster point identified by the counting model processes at least one viewpoint of the agent. The vector connecting the viewpoints and target points firmly indicates navigable space, since the propagation of light naturally points to unobstructed area. Besides, the vector can ensure that the crowd is in the FOV as well, since the target image is identified during ATE. Based on this prior, the ATE viewpoints are used to generate the final navigation points. The ATE view vector is calculated as:
@@
where $\eta$ is a hyper parameter representing the distance between the agent and the crowd cluster. NLBN ensures close-range, precise target observation and safe navigation, even in complex and unfamiliar environments. The potential navigation vectors, normal lines, and ATE view vectors are illustrated in Figure 4.
Figure 4: Illustration of the potential navigation vectors, normal lines, and FBE view vectors. Zoom in for better visualization.
@@ 3.3.4 Fine Detection and Counting (FDC)
Using the navigation points generated by NBLN, the agent travels from one point to another through path planning algorithms. Upon reaching each navigation point, close-range and high-resolution RGB observations can be conducted. These observations are then utilized as inputs for detection models to perform precise target detection. The detection results are projected onto the global point cloud using the depth sensor and the agent pose. To prevent repetitive detections, only one target is retained for each region within a specific scale. Ultimately, the result is calculated by counting the quantity of filtered targets.

### v2

@@ 3.3.1 Overview
Previous embodied navigation agents are designed for indoor environments long2024instructnav , or relying on language assistance liu2023aerialvln ; wang2024towards , making them restricted in large-scale outdoor scenes. Under such a context, a zero-shot agent, ZECC, which can actively control altitude and conduct crowd analysis, is proposed. As illustrated in Figure 3, the method consists of three components: Active Top-down Exploration (ATE), Normal-line based Navigation (NLBN), and Fine detection. ATE is designed for adjusting agent altitude for efficient coarse crowd distribution estimation, and NLBN estimates normal lines on top of the crowd for accurate crowd observation, alleviating occlusion.
@@ 3.3.2 Active Top-down Exploration (ATE)
Crowds are often concentrated in specific areas, such as roads and squares, making it unnecessary to explore every region of the environment. To address this, ATE is proposed to estimate global crowd distribution by changing altitude. This approach aims to improve efficiency by focusing exploration efforts on the most relevant areas. In particular, high-altitude exploration (HAE) brings a broader field of view for decision making, as well as relatively less exploration cost since obstacles are often sparse in high altitude. In contrast, low-altitude exploration (LAE) provides close-range observation for precise target detection. The agent needs to plan HAE and LAE to achieve efficiency and accuracy simultaneously. Therefore, ATE leverages the Z-axis mobility of outdoor agents and the common sense of MLLM to switch between HAE and LAE using local environment layout reasoning. Then, the crowd distribution is estimated by a crowd counting model to predict density maps.
Specifically, the agent collects observations ${O_{t}}=\left\{{o_{t}^{1},\ldots,o_{t}^{c}}\right\}$ and pose $p_{t}$ at time step $t$, where $c$ indicates the camera number of the agent. Then, an MLLM is prompted using observations and text prompts to conduct environment layout reasoning. The MLLM is asked to predict whether the current location is valuable for LAE, by referring to the current crowd appearance and obstacle layout. This process is formulated as:
@@
where $\text{MLLM}\left(\cdot\right)$ is the inference process and $I$ is the prompt. $s_{t}\in\left[{0,1}\right]$. If $s_{t}>0.5$, the agent will adjust its altitude for LAE. After switching the altitude strategy, the agent will conduct regular exploration. For HAE, one of the frontiers between explored and unknown areas is selected as a navigation point. For LAE, the agent keeps exploring until areas within its field of view during HAE are fully explored. During LAE, once the agent reaches a navigation site $f$, it gathers observations ${O_{f}}=\left\{{o_{f}^{1},\ldots,o_{f}^{c}}\right\}$, and a crowd counting model predicts crowd density maps on the observations. Then, the density maps are projected onto the global point cloud to form a global crowd distribution:
@@ 3.3.3 Normal-line based Navigation (NLBN)
NLBN is designed to analyze overlapping structures in dense crowds by converting the crowd detection task into surface detection, which helps identify optimal observation points. While random viewpoints may impair visibility, vantage points located above the center of the crowd provide clearer views for individual identification. This top-down perspective enables the distinction of overlapping individuals and maintains targets within the field of view (FOV). Initially, NLBN clusters the crowd point cloud into subregions using a clustering method, simplifying the analysis by breaking the large point cloud into manageable parts, as navigating to each point can be resource-intensive. Surfaces are then fitted to derive normal lines, thus transforming complex crowd analysis into a more straightforward surface analysis. Finally, optimized navigation points are generated based on these normal lines, ensuring they are elevated to facilitate accurate crowd detection while employing a viewpoint-based approach to avoid overlaps with obstacles.
In particular, GMM scrucca2023introduction is used to divide the global crowd distribution into manageable subregions. It can divide any crowd distribution into patches. A parameter $\epsilon$ is used to determine the size of each GMM cluster. Surfaces are fitted for each patch afterwards. Then, for the $N$-th cluster, the normal is obtained and represented by ${\bf{d}}_{N}^{{\rm{cluster}}}$. Candidate view directions $\{\mathbf{d}_{N1}^{\text{view}},...,\mathbf{d}_{Nm}^{\text{view}}\}_{N}$ are sampled under angular constraints:
@@
where $\zeta$ is a hyper parameter. These candidate view directions can bring the agent to optimized observation points by selecting a position along the vector. However, there are two issues: 1) The position may be located on obstacles; 2) The crowd cluster may be out of the agent’s FOV. Since the propagation of light naturally points to unobstructed areas, the ATE viewpoints are used to generate the final navigation points, which are calculated as:
@@
where $\eta$ is a hyper parameter representing the distance between the agent and the crowd cluster. NLBN ensures close-range, precise target observation and safe navigation, even in complex and unfamiliar environments. The potential navigation vectors, normal lines, and ATE view vectors are illustrated in Figure 2 (b).
@@ 3.3.4 Fine Detection and Counting (FDC)
Using the navigation points generated by NBLN, the agent travels from one point to another through path planning algorithms. Close-range and high-resolution RGB observations can be conducted upon reaching each navigation point. These observations are then utilized for detection models to perform precise target detection. The detection results are projected onto the global point cloud using the depth sensor and the agent’s pose. To prevent repetitive detections, only one target is retained for each region within a specific scale. Ultimately, the result is calculated by counting the number of filtered targets.